# RGI-Net: 3D Room Geometry Inference from Room Impulse Responses in the Absence of First-order Echoes

### v2

@@ Abstract
Room geometry is important prior information for implementing realistic 3D audio rendering. For this reason, various room geometry inference (RGI) methods have been developed by utilizing the time of arrival (TOA) or time difference of arrival (TDOA) information in room impulse responses. However, the conventional RGI technique poses several assumptions, such as convex room shapes, the number of walls known in priori, and the visibility of first-order reflections. In this work, we introduce the deep neural network (DNN), RGI-Net, which can estimate room geometries without the aforementioned assumptions. RGI-Net learns and exploits complex relationships between high-order reflections in room impulse responses (RIRs) and, thus, can estimate room shapes even when the shape is non-convex or first-order reflections are missing in the RIRs. The network takes RIRs measured from a compact audio device equipped with a circular microphone array and a single loudspeaker, which greatly improves its practical applicability. RGI-Net includes the evaluation network that separately evaluates the presence probability of walls, so the geometry inference is possible without prior knowledge of the number of walls.
@@ 1 Introduction
Sound propagation altered by room geometry is captured in RIRs in the form of various features. One of the representative features is the TOA identifiable from the propagation delay of its distinct peaks. Therefore, a lot of RGI studies have utilized TOA information [4, 5, 6, 7, 8, 9]. In particular, Antonacci et al. [4] localized 2D reflectors by converting a set of TOAs into ellipses. Two focal points of an ellipse correspond to a sound source and microphone, and the common tangent line across multiple ellipses can represent the room boundary. Baba et al. [8] extended ellipse-based method using stack-line detection in stacked RIRs. Common reflection lines were translated into real- and image-microphone positions, which determine room boundary. Lovedee-Turner and Murphy [9] demonstrated that the RGI of complex rooms, e.g., non-convex rooms, is possible based on RIRs acquired at multiple positions. They estimated all candidate walls and then filtered the estimated walls using a three-step validation process: path validation, line-of-sight (LOS) boundary validation, and closed geometry validation. The remaining walls constitute the final room geometry. However, the microphone array must be placed in a restricted area to secure the LOS condition for all the walls, and the acquired RIRs must include first-order reflections from the walls. Despite their success in estimating complex-shaped rooms, positioning sound sources at multiple locations makes it less practical.
@@
Most previous studies assume that first-order reflections from every wall are visible. For rooms with complex shapes, however, first-order reflection visibility cannot be ensured in one measurement with a compact microphone array. The fundamental remedy might be to utilize high-order reflections and analyze the complex relationship between low- and high-order reflections. Therefore, recent studies have demonstrated the potential of DNNs for high-level feature extraction from RIR [10, 11, 12, 13, 14, 15]. Yu and Kleijn [10], Poschadel et al. [11], and Tuna et al.[12] designed a DNN to extract the complex relationship between the temporal peaks of an RIR and estimated the geometrical parameters of a room. However, these DNN-based RGI studies have a limitation in that they cannot handle changes in the number of walls caused by various room types, as they have focused only on shoebox rooms. In our previous study [13], we attempted to address this limitation. However, the study only considered the LOS condition where all walls are visible from the audio device. Moreover, a spherical microphone array with 32 microphone capsules was used as the audio device, which makes the approach less practical.
@@ 3.1 RGI-Net Architecture
Fig. 2: Top view of rooms reconstructed from estimated wall parameters. The black dot (left) displays the position of a compact audio device. The black dashed lines and blue solid lines (right) correspond to reconstructed walls from the GT and inferred wall parameters, respectively.
@@ 3.2 Loss Function
In all the experiments conducted, we set $W_{0}\geq W$, and ground truth (GT) wall parameters and presence probabilities for nonexistent walls were initialized with zeroes. As the loss function for measuring the similarity between GT and estimated wall, angular loss between two flattened parameter vectors $\hat{{\mathbf{h}}}=\mathrm{flatten}(\hat{{\mathbf{A}}})$ and ${{\mathbf{h}}}=\mathrm{flatten}({{\mathbf{A}}})\in\mathbb{R}^{4W_{0}}$ was employed. That is,
@@
Along with the angular loss, we used a decision loss $L_{d}$ involved with the wall presence probability. The decision loss is defined as the binary cross entropy (BCE) between the GT and estimated probabilities: $p_{w}$ and $\hat{p}_{w}$. To cope with the order mismatch in the GT and estimated walls, the permutation invariant training technique [18] was employed during training. For training, the Adam [19] optimizer and cosine annealing learning rate scheduler [20] were used with the maximum learning rate of $1\text{e}^{-3}$.
@@ 4.1 Dataset
Although various measured RIRs have been released, their microphone-speaker configurations are not compatible with the compact audio device considered in this study. To train the model using RIRs obtainable from a compact audio system, we constructed an RIR dataset simulated using a circular microphone array with six omnidirectional microphones arranged on a circle of 0.05 m radius and a single loudspeaker centered in the array. The device was randomly positioned between $[1,~{}2]$ m from the floor and within 70% space of a given floorplan. In the simulated rooms, the floor and ceiling were parallel to each other and perpendicular to the other walls. Four different room types were considered: shoebox, pentagonal, hexagonal, and L-type rooms. Unlike the other types, L-type rooms can be categorized into L-LOS and L-NLOS (non-line-of-sight) types, depending on whether the device can capture LOS from all walls or not. The rooms were horizontally rotated within the range of $[0,~{}360]^{\circ}$ to reflect possible angular rotations of the device.
@@
The RIRs were simulated by ‘Pyroomacoustics’ [21] using the image source method (ISM) at a sampling rate of 8 kHz. The absorption coefficient of each room was randomly selected within $[0.1,~{}0.3]$. The distance between the loudspeaker and microphones was always kept constant, so we trimmed out the direct part and silence region corresponding to the initial propagation delay from RIRs. Two datasets with low- and high-noise levels were constructed by adding white Gaussian noise to RIRs. In the high-noise level dataset, the noise level was sufficiently high to mask the peaks from third- or higher-order reflections. The resultant train dataset contains 600k RIRs, including 1k validation data.
@@
Table 2: Performance comparison to different RGI methods.
4.2 Evaluation Metrics
4.3 Experimental Results
The RGI results on two different noise levels are listed in Table 1. The overall result shows that only the existence of 2 and 11 walls out of 4k candidate walls was misclassified in the low- and high-noise datasets, respectively. The average errors are approximately 10 cm and ${}^{\circ}$ for the low-noise dataset. RGI-Net shows good performance for most room types, but the most significant error variation between low- and high-noise levels is observed in L-NLOS rooms. Fig. 2 depicts the comparison between GT and inferred rooms based on the high-noise dataset. The left part of each result shows the location of the audio device (black dot), which was set as the origin of coordinates during inference. These results show that most of the inferred walls (blue solid lines) are very close to the GT walls (black dashed lines), while the invisible wall shown in Fig. 2(d) exhibits a more significant error than the other walls in the L-NLOS room. This is because the strong background noise masks high-order reflections that are small in magnitude but crucial in estimating the invisible walls in the L-NLOS rooms. This indirectly indicates that RGI-Net utilizes higher-order reflections when some low-order reflections are missing.
@@ 4.3 Experimental Results
Next, we compare the proposed model with different RGI techniques. Table 2 describes the performance comparison with two conventional [8, 9] and one DNN-based [12] methods. Although direct comparison is not possible due to the differences in their data and conditions, the errors of the proposed method shown in Table 2 are in a similar range to those of the conventional and DNN-based techniques under low-noise conditions. More importantly, the proposed model can estimate non-convex rooms without relocation of the audio device to secure the LOS condition, and without prior knowledge of the number of walls.

### v4

# RGI-Net: 3D Room Geometry Inference from Room Impulse Responses With Hidden First-Order Reflections

@@ Abstract
Room geometry is important prior information for implementing realistic 3D audio rendering. For this reason, various room geometry inference (RGI) methods have been developed by utilizing the time-of-arrival (TOA) or time-difference-of-arrival (TDOA) information in room impulse responses (RIRs). However, the conventional RGI technique poses several assumptions, such as convex room shapes, the number of walls known in priori, and the visibility of first-order reflections. In this work, we introduce the RGI-Net which can estimate room geometries without the aforementioned assumptions. RGI-Net learns and exploits complex relationships between low-order and high-order reflections in RIRs and, thus, can estimate room shapes even when the shape is non-convex or first-order reflections are missing in the RIRs. RGI-Net includes the evaluation network that separately evaluates the presence probability of walls, so the geometry inference is possible without prior knowledge of the number of walls.
@@ 1 Introduction
Sound propagation altered by room geometry is captured in room impulse responses (RIRs) in the form of various features. One of the representative features is the time-of-arrival (TOA) identifiable from the propagation delay of its distinct peaks. Therefore, a lot of room geometry inference (RGI) studies have utilized TOA information [4, 5, 6, 7, 8, 9]. In particular, Antonacci et al. [4] localized 2D reflectors by converting a set of TOAs into ellipses. Two focal points of an ellipse correspond to a sound source and microphone, and the common tangent line across multiple ellipses can represent the room boundary. Baba et al. [8] extended ellipse-based method using stack-line detection in stacked RIRs. Common reflection lines were translated into real- and image-microphone positions, which determine room boundary. Lovedee-Turner and Murphy [9] demonstrated that the RGI of complex rooms, e.g., non-convex rooms, is possible based on RIRs acquired at multiple positions. They estimated all candidate walls and then filtered the estimated walls using a three-step validation process: path validation, line-of-sight (LOS) boundary validation, and closed geometry validation. The remaining walls constitute the final room geometry. However, the microphone array must be placed in a restricted area to secure the LOS condition for all the walls, and the acquired RIRs must include first-order reflections from the walls. Despite their success in estimating complex-shaped rooms, positioning sound sources at multiple locations makes it less practical.
@@
Most previous studies assume that first-order reflections from every wall are visible. For rooms with complex shapes, however, first-order reflection visibility cannot be ensured in one measurement with a compact microphone array. The fundamental remedy might be to utilize high-order reflections and analyze the complex relationship between low- and high-order reflections. Therefore, recent studies have demonstrated the potential of deep neural networks (DNNs) for high-level feature extraction from RIR [10, 11, 12, 13, 14, 15]. Yu and Kleijn [10], Poschadel et al. [11], and Tuna et al.[12] designed a DNN to extract the complex relationship between the temporal peaks of an RIR and estimated the geometrical parameters of a room. However, these DNN-based RGI studies have a limitation in that they cannot handle changes in the number of walls caused by various room types, as they have focused only on shoebox rooms. In our previous study [13], we attempted to address this limitation. However, the study only considered the LOS condition where all walls are visible from the audio device. Moreover, a spherical microphone array with 32 microphone capsules was used as the audio device, which makes the approach less practical.
@@ 3.2 Loss Function
In all the experiments conducted, we set $W_{0}=8\geq W$, and ground truth (GT) wall parameters and presence probabilities for nonexistent walls were initialized with zeroes. As the loss function for measuring the similarity between GT and estimated wall, angular loss between two flattened parameter vectors $\hat{{\mathbf{h}}}=\mathrm{flatten}(\hat{{\mathbf{A}}})$ and ${{\mathbf{h}}}=\mathrm{flatten}({{\mathbf{A}}})\in\mathbb{R}^{4W_{0}}$ was employed. That is,
@@
Along with the angular loss, we used a decision loss $L_{d}$ involved with the wall presence probability. The decision loss is defined as the binary cross entropy (BCE) between the GT and estimated probabilities: $p_{w}$ and $\hat{p}_{w}$. To cope with the order mismatch in the GT and estimated walls, the permutation invariant training technique [18] was employed during training. For training, the Adam [19] optimizer and cosine annealing learning rate scheduler [20] were used with the maximum learning rate of $10^{-3}$.
@@ 4.1 Dataset
Although various measured RIRs have been released, their microphone-speaker configurations are not compatible with the compact audio device considered in this study. To train the model using RIRs obtainable from a compact audio system, we constructed an RIR dataset simulated using a circular microphone array with six omnidirectional microphones arranged on a circle of 0.05 m radius and a single loudspeaker centered in the array. The device was randomly positioned between $[1,~{}2]$ m from the floor and within 70% space defined by equally scaling down from every vertex of a given floorplan of the room. In the simulated rooms, the floor and ceiling were parallel to each other and perpendicular to the other walls. Four different room types were considered: shoebox, pentagonal, hexagonal, and L-type rooms. Unlike the other types, L-type rooms can be categorized into L-LOS and L-NLOS (non-line-of-sight) types, depending on whether the device can capture LOS from all walls or not. The rooms were horizontally rotated within the range of $[0,~{}360]^{\circ}$ to reflect possible angular rotations of the device.
@@
The RIRs were simulated by ‘Pyroomacoustics’ [21] using the image source method (ISM) at a sampling rate of 8 kHz. The absorption coefficient of each room was randomly selected within $[0.1,~{}0.3]$. The loudspeaker and microphones kept a consistent distance, so we trimmed out the direct part of the RIRs. Two datasets with low- and high-noise levels were constructed by adding white Gaussian noise to RIRs. The noise level was adjusted to maintain a signal-to-noise ratio (SNR) within the ranges of $[20,~{}30]$ and $[10,~{}20]$ dB for low- and high-noise level datasets, respectively. In the high-noise level dataset, the noise level was sufficiently high to mask the peaks from third- or higher-order reflections. The resultant train dataset contains 600k RIRs, including 1k validation data.
Fig. 2: Top view of rooms reconstructed from estimated wall parameters. Since four distinct L-shaped rooms can be formed by the estimated planes, (c) and (d) were reconstructed considering the GT room shapes. The black dot (left) denotes the position of an audio device. The black dashed lines and blue solid lines (right) correspond to reconstructed walls from the GT and inferred wall parameters, respectively.
4.2 Evaluation Metrics
4.3 Experimental Results
The RGI results on two different noise levels are listed in Table 1. The overall result shows that only the existence of 2 and 11 walls out of 4k candidate walls was misclassified in the low- and high-noise datasets, respectively. Therefore, the $ACC_{W}$ of the low- and high-noise datasets are 99.95% and 99.72%, respectively. The average errors are approximately 10 cm and 1.9 for the low-noise dataset. RGI-Net shows good performance for most room types, but the most significant error variation between low- and high-noise levels is observed in L-NLOS rooms. Fig. 2 depicts the comparison between GT and inferred rooms based on the high-noise dataset. The left part of each result shows the location of the audio device (black dot), which was set as the origin of coordinates during inference. These results show that most of the inferred walls (blue solid lines) are very close to the GT walls (black dashed lines), while the invisible wall shown in Fig. 2(d) exhibits a more significant error than the other walls in the L-NLOS room. This is because the strong background noise masks high-order reflections that are small in magnitude but crucial in estimating the invisible walls in the L-NLOS rooms. This indirectly indicates that RGI-Net utilizes higher-order reflections when some low-order reflections are missing.
Next, we describe the RGI performance reported from two conventional [8, 9] and one DNN-based [12] methods. The errors of the proposed method shown in Table 2 are in a similar range to those of the conventional and DNN-based techniques under low-noise conditions. However, this table is not for direct comparison across the methods since there are differences in their room setup, source-microphone configurations, and assumptions. More importantly, the inference of non-convex rooms without relocation of the audio device to secure the LOS condition, and without prior knowledge of the number of walls is a key ability of RGI-Net.
Table 2: Performance comparison to different RGI methods.